# Rethinking of Feature Interaction for Multi-task Learning on Dense Prediction

### v1

@@ 4.2.1 Overall Comparisons
We compared BFCI with various state-of-the-art multi-task methods on the NYUD-v2 and PASCAL Context datasets. We employed InternImage-L as the backbone. As shown in Table I, across all the compared tasks, our method surpasses previous works significantly in all of tasks. On the NYUD-v2 dataset, the tasks include semantic segmentation (Seg.), depth estimation (Dep.), surface normal estimation (Norm.), and edge detection (Edge.). For the NYUD-v2 comparisons, we used TFR-huge, and BFCI outperforms previous methods on all tasks. Particularly significant improvements are achieved in semantic segmentation and depth estimation. Compared to InvPT [50], our method enhances by $+3.01$ mIoU and $-0.0528$ rmse, respectively. Compared to the previous state-of-the-art multi-scale interaction MTI-Net [38], we achieve a significant improvement of $+10.60$ mIoU and $-0.0710$ rmse. This improvement is substantial. Noticeable enhancements are also seen in the other two tasks. Qualitative comparisons are shown in Fig. 6, it’s evident that our method yields better prediction quality for each task.
@@
On the PASCAL Context dataset, the tasks include semantic segmentation (Seg.), human body part segmentation (H.Parts), surface normal estimation (Norm.), saliency estimation (Sal.), and edge detection (Edge.). We used TFR-large for comparison. Due to the increased number of tasks, achieving a balance in performance across multiple tasks becomes challenging, and maintaining superiority across all tasks over previous works is not easy to achieve. Nevertheless, even under these conditions, we achieved balanced promotions in all tasks. More significant improvement occurred in human body part segmentation (H.Parts) and saliency estimation (Sal.), resulting in $+3.72$ mIoU and $-0.77$ mErr. Besides, we also conduct qualitative comparisons in Fig. 7. Furthermore, we visualized the task-specific features refined through TFR, as shown in Fig. 8. In comparison to works lacking the TFR structure, such as MTI-Net [38], the task-specific features generated by our method exhibit clearer high-level semantics, highlighted regions possess sharper boundaries, and overall coherence is improved.
@@
Figure 8: The visualization of feature maps at scale 0 by MTI-Net (a) and our BFCI (b). In semantic segmentation (left), our method presents more consistency with clear edges for the same semantic region. The brighter area means more attention is paid in those areas. In depth estimation (right), our method also produces more accurate edges for areas with different depths, such as the door in the faraway distance which is supposed to be the brightest area. $T_{1}$ and $T_{2}$ represent semantic and depth labels respectivelt.
@@ 4.2.2 Comparisons on Different Backbones
Different from InvPT [50] which is specially designed for ViT backbones, our method is adaptive to different types of backbones. We also employed different backbones to compare with various established methods, ensuring the effectiveness of our method across diverse conditions. We utilized two different sizes of ViTs [10], ViT-base and ViT-large, two different sizes of HRNet [35], HRNet-w18 and HRNet-w48, as well as the commonly used ResNet-50d with dilated convolutions [16, 6]. As observed, our BFCI demonstrates robust performance across various types of backbone networks.
@@
When using the HRNet series backbone networks, our method exhibits comparable performance to other methods in the Semseg. task, while showcasing significant improvements in other spatial geometric estimation tasks. On ResNet-50d, though our method achieves competitive performance, the performance among different tasks is more balanced, which means our method does not tend to sacrifice performance on one task to enhance another. When employing larger parameter backbone networks such as the ViT series, our method shows significant enhancements in high-level semantic understanding tasks like Semseg. and H.Parts. The reason behind this phenomenon is that when the backbone network has fewer parameters, its feature extraction capacity might be limited. Simpler geometric estimation tasks (Norm., Edge.) might not have reached saturation, allowing the model to learn useful representations from richer semantic tasks (Semseg.) and thereby enhance the performance. Conversely, when the backbone network has more parameters, it can capture precise image representations, reaching saturation in geometric estimation tasks. At this point, further improvements become challenging. However, a larger feature space and richer representations become advantageous for tasks that require advanced information (Semseg. and H.Parts). This phenomenon is also observed in [50].
4.2.3 Model Size Comparisons
@@
To further show the advantages of our BFCI, we compare our method with some of the previous SOTA works with the consideration of computation cost and model parameter amounts. As shown in Table III, under the same ViT-large backbones, out method shows clear superiority over previous works on the majority of tasks, especially on Seg., Sal. and Norm. Benefits from the high-efficiency interactions brought by BFE, our BFCI reduces 24 GFLOPs than InvPT, even with the configuration of TFR-large. For different backbone configurations, our method shows better performance with less computation costs on InternImage backbones. Compared with ViT-base, our method performs better with compatible FLOPs and Params on InternImage-B. When compared with ViT-large, our method achieves better performance on InternImage-L even with fewer parameters (-57 MParams) and less computation costs (-118 GFLOPs).
@@
TABLE IV: Ablation analysis of the components of BFCI. The experiments are conducted on NYUD-v2.
@@
TABLE V: Ablation analysis of refine strategies. The experiments are conducted on NYUD-v2.
@@
TABLE VI: Comparison of different task sets on NYUD-v2.
@@
TABLE VII: Comparison of different task sets, we use the TFR-base in decoding stage.
@@ 4.3 Ablation Study
The ablation experiments are divided into three sections. In section 4.3.1, Model Analysis, we delve into the significance of the three core modules that constitute BFCI: BFE, TPP, and TFR. We provide comprehensive visual analysis in this regard. In section 4.3.2, Refine Strategies Analysis, we explore how different methods of feature refinement impact the final predictive performance, including varying sizes of TFR and other techniques. In section 4.3.3, Task Set Analysis, we investigate how the model’s performance changes when dealing with different combinations of tasks in multi-task learning. For all the experiments, we employ ViT-base as the backbone network.
@@ 4.3.1 Model Analysis
These experiments were conducted on the NYUD-v2 dataset. Initially, we trained and tested the performance of single-task learning (STL) on each individual task. Then, we calculated the performance of the multi-task learning baseline (MTL baseline) and progressively incorporated BFE, TPP, and TFR. The comparative results are presented in Table IV. As evident, due to the varying number of tasks and their differences, each task pursues a distinct optimization objective. Consequently, MTL baseline exhibits a noticeable performance decline in comparison to STL. Upon introducing BFE over the baseline, performance improvement is observed, nearing the level of advanced models like InvPT [50] shown in Table II. However, at this stage, without TPP’s assistance in obtaining more discriminative task-specific features, the quality of the Bridge Features generated by BFE is not optimal. Hence, with the introduction of TPP, further performance enhancement is achieved.
@@
We also conducted comprehensive visual analysis. As shown in Fig. 9. We selected bridge features ($\mathbf{S^{\prime}}$) and compared them with task-generic features ($\mathbf{S}$) and task-specific features ($\mathbf{P}$) at different scales. These results are presented in Fig. 9. The distribution of task-specific features $\mathbf{P}$ is similar to their corresponding labels and tends to emphasize the relevant regions under supervision. For example, they highlight regions with significant depth variations, as well as distinct semantic areas like floors, walls, and cabinets. This suggests that they possess distinct task-specific perception capabilities. On the other hand, task-generic features $\mathbf{S}$ exhibit rich low-level representations, such as edges and textures. Bridge features $\mathbf{S^{\prime}}$ exhibit both task-specific perception capabilities and low-level representations. This demonstrates that $\mathbf{S^{\prime}}$ has the potential to serve as an intermediate medium for multi-task interaction, ensuring the integrity of feature interactions.
@@ 4.3.3 Task Set Analysis
Table VI illustrates different combinations of tasks on the NYUD-v2 dataset. We considered 6 distinct task sets, including cases with 2, 3, and 4 tasks. It can be observed that when fewer tasks are involved, the occurrence of mutual enhancement between tasks becomes more likely. Notably, the most significant mutual enhancement happens between the Seg. and Dep. tasks (Seg. $+3.49\%$ $\Delta_{\tau_{seg}}$, Dep. $+7.92\%$ $\Delta_{\tau_{dep}}$), resulting in an average performance improvement of $+5.70\%\Delta_{MTL}$. This aligns with findings from several studies, indicating a strong latent correlation between depth estimation and semantic segmentation. However, the correlation between the Seg. and Norm. tasks is relatively weak, leading to considerable negative transfer effects (Seg. $-1.28\%$ $\Delta_{\tau_{seg}}$, Norm. $+0.26\%$ $\Delta_{\tau_{norm}}$), resulting in an average performance decline of $-0.51\%\Delta_{MTL}$. The surface normals are aligned with the gradients of depth in the space, establishing a strong correlation between these two tasks, facilitating mutual enhancement (Dep. $+4.95\%$ $\Delta_{\tau_{dep}}$, Norm. $+0.94\%$ $\Delta_{\tau_{norm}}$), and achieving an average performance improvement of $+2.94\%\Delta_{MTL}$. As the number of tasks increases, due to inherent inconsistencies between various task pairs, such as Seg. and Norm., performance tends to decrease. For example, the combination of Seg., Dep., and Norm. results in $+1.90\%\Delta_{MTL}$, with a more noticeable decline in Seg.’s performance. In contrast, the combination of Dep., Norm., and Edge. exhibits good performance, with an average improvement of $+2.07\%\Delta_{MTL}$, as all three tasks are designed to learn spatial geometry and lack evident conflicts.
@@
Table VII depicts different combinations of tasks on the PASCAL Context dataset. We also examined 6 different task combinations, encompassing 2, 3, 4, and 5 tasks. Among these, the Seg., H.Parts, and Sal. tasks exhibit higher correlation. Learning these tasks together or in pairs does not yield noticeable negative effects. Seg. and H.Parts are tasks involving segmentation at different granularities, and they share a substantial amount of information. This direct information sharing explicitly enhances prediction accuracy of the person category in Seg., boosting the Intersection over Union (IoU) from $86.10$ to $86.67$. The interaction of high-level semantic information across different granularities benefits both tasks, resulting in an average performance improvement of $+1.20\%\Delta_{MTL}$. Similarly, the Seg. and Sal. tasks share some foreground and background information, leading to no substantial negative impact, with an average performance change of $-0.22\%\Delta_{MTL}$. Learning all three tasks together results in an average performance change of $-0.33\%\Delta_{MTL}$, which is an acceptable compromise. However, introducing Norm. in addition to these three tasks significantly reduces the gains, with an average performance decline of $-2.70\%\Delta_{MTL}$. This observation aligns with the findings on the NYUD-v2 dataset. Furthermore, the addition of Edge. results in a performance rebound, enhancing the relative gains for nearly every task. This underscores the importance of Edge. for most dense prediction tasks. When learning Sal., Norm., and Edge., the three relatively simpler tasks together, the average performance is $+0.31\%\Delta_{MTL}$.
@@ 5 Conclusion
In this paper, we discovered the limitations of previous multi-task dense prediction methods in terms of information loss between tasks and inadequate exploration of multi-task feature interactions. To address these challenges, we introduced a novel approach called Bridge-Feature-Centric Interaction (BFCI). Specifically, our method employs a shared encoder backbone to generate task-generic features. In the early decoding stage, a preliminary decoder with Task Pattern Propagation (TPP) is used to generate high-quality task-specific features. We introduced the Bridge Feature Extractor (BFE) to interact between globally queried task-specific features and backbone task-generic features, selecting task-specific perception information and generating multi-scale bridge features. The Task Feature Refiner (TFR) injects the representation of bridge features and iteratively optimizes the final task predictions, resulting in the ultimate task predictions. Compared to previous structures, our proposed method ensures the integrity of feature interaction, striking a good balance between performance and cost.
@@
In our experiments, we evaluated our method on the widely used multi-task learning benchmark datasets, NYUD-v2 and PASCAL Context. By comparing with Single Task Learning (STL), Multi-Task Learning Baseline (MTL baseline), PAD-Net, MTI-Net, InvPT, and our BFCI, we observed the superior performance of BFCI across multiple tasks. Furthermore, we conducted comprehensive visual analysis to demonstrate the advantage of bridge features over the other two types (task-generic and task-specific features). Compared to models generating and utilizing task-generic and task-specific features, such as BMTAS and InvPT, BFCI dives deeper into feature mining from important dense prediction priors, driving significant performance improvements. We analyzed the reasons for our method’s superiority at the feature level. Moreover, through ablation experiments on model components, we analyzed the effectiveness of each component and demonstrated the improvements of our method tailored for dense prediction tasks. Additionally, our ablation experiments on task sets started from the nature of each task to analyze task relationships and studied the impact of different task combinations on multi-task learning performance. By conducting these ablation experiments on task sets, we gained a better understanding of the interactions between tasks and provided guidance for constructing effective multi-task learning models.
@@
In summary, our research addressed the limitations of previous multi-task dense prediction methods through the introduction of the BFCI approach. Experimental results show that BFCI achieves significant performance improvements in multi-task learning, which were further validated through visual analysis and ablation experiments. Our work provides valuable insights and guidance for research and practice in the field of multi-task dense prediction. Future research can further explore the application and optimization of the BFCI method to meet the broader demands of dense prediction tasks.

### v2

@@ 4.2.1 Overall Comparisons
We compared BridgeNet with various state-of-the-art multi-task methods on the NYUD-v2, Cityscapes and PASCAL Context datasets. We employed InternImage-L as the backbone for training on NYUD-v2 and PASCAL Context, and ViT-large as the backbone for Cityscapes. As shown in Table I, across all the compared tasks, our method surpasses previous works significantly in all of the tasks. On the NYUD-v2 dataset, the tasks include semantic segmentation (Seg.), depth estimation (Dep.), surface normal estimation (Norm.), and edge detection (Edge.). For the NYUD-v2 comparisons, we used TFR-huge, and BridgeNet outperforms previous methods on all tasks. Particularly significant improvements are achieved in semantic segmentation and depth estimation. Compared to InvPT [51], our method enhances by $+3.01$ mIoU and $-0.0528$ rmse, respectively. Compared to the previous state-of-the-art multi-scale interaction MTI-Net [39], we achieve a significant improvement of $+10.60$ mIoU and $-0.0710$ rmse. This improvement is substantial. Noticeable enhancements are also seen in the other two tasks. Qualitative comparisons are shown in Fig. 6, which further prove that BridgeNet yields better prediction quality.
On the Cityscapes, the tasks include semantic segmentation (Seg.) and depth estimation (Dep.), and we use TFR-huge. Compared with encoder-focused methods which only utilize low-level representations in task-generic features for task interactions, our method significantly achieves better performance on both tasks. Also, our BridgeNet surpasses the best performing previous decoder-focused InvPT [51] by $+1.42$ mIoU and $-0.025$ rmse. We also make qualitative comparisons in Fig. 7 to further validate the effectiveness of our method.
@@
On the PASCAL Context dataset, the tasks include semantic segmentation (Seg.), human body part segmentation (H.Parts), surface normal estimation (Norm.), saliency estimation (Sal.), and edge detection (Edge.). We used TFR-large for BridgeNet. Due to the increased number of tasks, achieving a balance in performance across multiple tasks becomes challenging, and maintaining superiority across all tasks over previous works is not easy to achieve. Nevertheless, even under these conditions, we achieved balanced promotions in all tasks. More significant improvement occurred in human body part segmentation (H.Parts) and saliency estimation (Sal.), resulting in $+3.72$ mIoU and $-0.77$ mErr. Besides, we also conduct qualitative comparisons in Fig. 8. Furthermore, we visualized the task-specific features refined through TFR, as shown in Fig. 9. In comparison to works lacking the TFR structure, such as MTI-Net [39], the task-specific features generated by our method exhibit clearer high-level semantics, highlighted regions possess sharper boundaries, and overall coherence is improved.
@@
Figure 9: The visualization of feature maps at scale 0 by MTI-Net (a) and our BridgeNet (b). In semantic segmentation (left), our method presents more consistency with clear edges for the same semantic region. The brighter area means more attention is paid in those areas. In depth estimation (right), our method also produces more accurate edges for areas with different depths, such as the door in the faraway distance which is supposed to be the brightest area. $T_{1}$ and $T_{2}$ represent semantic segmentation and depth estimation labels respectively.
@@ 4.2.2 Comparisons on Different Backbones
Different from InvPT [51] which is specially designed for ViT backbones, our method is adaptive to different types of backbones. We also employed different backbones to compare with various established methods, ensuring the effectiveness of our method across diverse conditions. We utilized two different sizes of ViTs [11], ViT-base and ViT-large, two different sizes of HRNet [36], HRNet-w18 and HRNet-w48, as well as the commonly used ResNet-50d with dilated convolutions [17, 6]. As observed, our BridgeNet demonstrates robust performance across various types of backbone networks.
@@
When using the HRNet series backbone networks, our method exhibits comparable performance to other methods in the Semseg. task, while showcasing significant improvements in other spatial geometric estimation tasks. On ResNet-50d, though our method achieves competitive performance, the performance among different tasks is more balanced, which means our method does not tend to sacrifice performance on one task to enhance another. When employing larger parameter backbone networks such as the ViT series, our method shows significant enhancements in high-level semantic understanding tasks like Semseg. and H.Parts. The reason behind this phenomenon is that when the backbone network has fewer parameters, its feature extraction capacity might be limited. Simpler geometric estimation tasks (Norm., Edge.) might not have reached saturation, allowing the model to learn useful representations from richer semantic tasks (Semseg.) and thereby enhance the performance. Conversely, when the backbone network has more parameters, it can capture precise image representations, reaching saturation in geometric estimation tasks. At this point, further improvements become challenging. However, larger feature space and richer representations become advantageous for tasks that require advanced information (Semseg. and H.Parts). This phenomenon is also observed in [51].
4.2.3 Comparisons on Model Size and Cost
@@
To further show the efficiency of our BridgeNet, we compare our method with some of the previous SOTA works with the consideration of computation cost, parameter amount, latency, and memory cost regarding a single image input. We test all models on a single NVIDIA 3090 GPU under the same condition. As shown in Table IV, our method achieves competitive performance with ViT-base compared with other methods applied with ViT-large, and under the same ViT-large backbones, our method shows clear superiority over previous works on the majority of tasks, especially on Seg., Sal. and Norm. Benefits from the high-efficiency interactions brought by BFE, our BridgeNet with the configuration of TFR-large reduces over 100 GFLOPs compared with [49, 39, 3], and 24 GFLOPs compared with InvPT [51]. Considering the memory cost for single image inference, our BridgeNet significantly surpasses methods with heavy Multi-scale Feature Propagation Module (FPM) like [39, 3], with using less than half of the memory. Though our method is slightly slower when comparing the inference latency with other methods, which is due to the two-step extraction and refinement involving the bridge feature, however the advantages in model size, FLOPs, memory consumption, etc. prove that our BridgeNet saves a lot of additional computational overhead for multi-task interactions.
For different backbone configurations, our method shows better performance with less computation costs on InternImage backbones. Compared with ViT-base, our method performs better with compatible FLOPs and Params on InternImage-B. When compared with ViT-large, our method achieves better performance on InternImage-L even with fewer parameters (-57 MParams) and less computation costs (-118 GFLOPs).
@@
TABLE V: Ablation analysis of the components of BridgeNet. The experiments are conducted on NYUD-v2.
@@
TABLE VI: Ablation analysis of refine strategies. The experiments are conducted on NYUD-v2.
@@
TABLE VII: Comparison of different task sets on NYUD-v2.
@@
TABLE VIII: Comparison of different task sets, we use the TFR-base in decoding stage.
@@ 4.3 Ablation Study
The ablation experiments are divided into three sections. In section 4.3.1, Model Analysis, we delve into the significance of the three core modules that constitute BridgeNet: BFE, TPP, and TFR. We provide comprehensive visual analysis in this regard. In section 4.3.2, Refine Strategies Analysis, we explore how different methods of feature refinement impact the final predictive performance, including varying sizes of TFR and other techniques. In section 4.3.3, Task Set Analysis, we investigate how the model’s performance changes when dealing with different combinations of tasks in multi-task learning. For all the experiments, we employ ViT-base as the backbone network.
@@ 4.3.1 Model Analysis
These experiments were conducted on the NYUD-v2 dataset. Initially, we trained and tested the performance of single-task learning (STL) on each individual task. Then, we calculated the performance of the multi-task learning baseline (MTL baseline) and progressively incorporated BFE, TPP, and TFR. The comparative results are presented in Table V. As evident, due to the varying number of tasks and their differences, each task pursues a distinct optimization objective. Consequently, MTL baseline exhibits a noticeable performance decline in comparison to STL. Upon introducing BFE over the baseline, performance improvement is observed, nearing the level of advanced models like InvPT [51] shown in Table III. However, at this stage, without TPP’s assistance in obtaining more discriminative task-specific features, the quality of the Bridge Features generated by BFE is not optimal. Hence, with the introduction of TPP, further performance enhancement is achieved.
@@
We also conducted a comprehensive visual analysis. As shown in Fig. 10. We selected bridge features ($\mathbf{S^{\prime}}$) and compared them with task-generic features ($\mathbf{S}$) and task-specific features ($\mathbf{P}$) at different scales. These results are presented in Fig. 10. The distribution of task-specific features $\mathbf{P}$ is similar to their corresponding labels and tends to emphasize the relevant regions under supervision. For example, they highlight regions with significant depth variations, as well as distinct semantic areas like floors, walls, and cabinets. This suggests that they possess distinct task-specific perception capabilities. On the other hand, task-generic features $\mathbf{S}$ exhibit rich low-level representations, such as edges and textures. Bridge features $\mathbf{S^{\prime}}$ exhibit both task-specific perception capabilities and low-level representations. This demonstrates that $\mathbf{S^{\prime}}$ has the potential to serve as an intermediate medium for multi-task interaction, ensuring the integrity of feature interactions.
@@ 4.3.3 Task Set Analysis
Table VII illustrates different combinations of tasks on the NYUD-v2 dataset. We considered 6 distinct task sets, including cases with 2, 3, and 4 tasks. It can be observed that when fewer tasks are involved, the occurrence of mutual enhancement between tasks becomes more likely. Notably, the most significant mutual enhancement happens between the Seg. and Dep. tasks (Seg. $+3.49\%$ $\Delta_{\tau_{seg}}$, Dep. $+7.92\%$ $\Delta_{\tau_{dep}}$), resulting in an average performance improvement of $+5.70\%\Delta_{MTL}$. This aligns with findings from several studies, indicating a strong latent correlation between depth estimation and semantic segmentation. However, the correlation between the Seg. and Norm. tasks are relatively weak, leading to considerable negative transfer effects (Seg. $-1.28\%$ $\Delta_{\tau_{seg}}$, Norm. $+0.26\%$ $\Delta_{\tau_{norm}}$), resulting in an average performance decline of $-0.51\%\Delta_{MTL}$. The surface normals are aligned with the gradients of depth in the space, establishing a strong correlation between these two tasks, facilitating mutual enhancement (Dep. $+4.95\%$ $\Delta_{\tau_{dep}}$, Norm. $+0.94\%$ $\Delta_{\tau_{norm}}$), and achieving an average performance improvement of $+2.94\%\Delta_{MTL}$. As the number of tasks increases, due to inherent inconsistencies between various task pairs, such as Seg. and Norm., performance tends to decrease. For example, the combination of Seg., Dep., and Norm. results in $+1.90\%\Delta_{MTL}$, with a more noticeable decline in Seg.’s performance. In contrast, the combination of Dep., Norm., and Edge. exhibits good performance, with an average improvement of $+2.07\%\Delta_{MTL}$, as all three tasks are designed to learn spatial geometry and lack evident conflicts.
@@
Table VIII depicts different combinations of tasks on the PASCAL Context dataset. We also examined 6 different task combinations, encompassing 2, 3, 4, and 5 tasks. Among these, the Seg., H.Parts, and Sal. tasks exhibit higher correlation. Learning these tasks together or in pairs does not yield noticeable negative effects. Seg. and H.Parts are tasks involving segmentation at different granularities, and they share a substantial amount of information. This direct information sharing explicitly enhances prediction accuracy of the person category in Seg., boosting the Intersection over Union (IoU) from $86.10$ to $86.67$. The interaction of high-level semantic information across different granularities benefits both tasks, resulting in an average performance improvement of $+1.20\%\Delta_{MTL}$. Similarly, the Seg. and Sal. tasks share some foreground and background information, leading to no substantial negative impact, with an average performance change of $-0.22\%\Delta_{MTL}$. Learning all three tasks together results in an average performance change of $-0.33\%\Delta_{MTL}$, which is an acceptable compromise. However, introducing Norm. in addition to these three tasks significantly reduces the gains, with an average performance decline of $-2.70\%\Delta_{MTL}$. This observation aligns with the findings on the NYUD-v2 dataset. Furthermore, the addition of Edge. results in a performance rebound, enhancing the relative gains for nearly every task. This underscores the importance of Edge. for most dense prediction tasks. When learning Sal., Norm., and Edge., the three relatively simpler tasks together, the average performance is $+0.31\%\Delta_{MTL}$.
@@ 5 Conclusion
In this paper, we discovered the limitations of previous multi-task dense prediction methods in terms of incomplete levels of representations, less discriminative semantics in feature participants, and inefficient pair-wise task interaction processes. To address these challenges, we introduced a novel framework called BridgeNet. Specifically, our method adopts a shared encoder backbone to generate task-generic features. In the early decoding stage, a preliminary decoder with Task Pattern Propagation (TPP) is used to generate high-quality task-specific features. We introduced the Bridge Feature Extractor (BFE) to interact between globally queried task-specific features and backbone task-generic features, selecting task-specific perception information and generating multi-scale bridge features. The Task Feature Refiner (TFR) injects the representation of bridge features and iteratively optimizes the final task predictions, resulting in the ultimate task predictions. Compared to previous structures, our proposed method ensures the integrity of feature interaction, striking a good balance between performance and cost.
@@
In our experiments, we evaluated our method on the widely used multi-task learning benchmark datasets, NYUD-v2, Cityscapes and PASCAL Context. By comparing with Single Task Learning (STL), Multi-Task Learning Baseline (MTL baseline), PAD-Net, MTI-Net, InvPT, and our BridgeNet, we observed the superior performance of BridgeNet across multiple tasks. Furthermore, we conducted a comprehensive visual analysis to demonstrate the advantage of bridge features over the other two types (task-generic and task-specific features). Compared to models generating and utilizing task-generic and task-specific features, such as BMTAS and InvPT, BridgeNet dives deeper into feature mining from important dense prediction priors, driving significant performance improvements. We analyzed the reasons for our method’s superiority at the feature level. Moreover, through ablation experiments on model components, we analyzed the effectiveness of each component and demonstrated the improvements of our method tailored for dense prediction tasks. Additionally, our ablation experiments on task sets started from the nature of each task to analyze task relationships and studied the impact of different task combinations on multi-task learning performance. By conducting these ablation experiments on task sets, we gained a better understanding of the interactions between tasks and provided guidance for constructing effective multi-task learning models.
@@
In summary, our research addressed the limitations of previous multi-task dense prediction methods through the introduction of the BridgeNet approach. Experimental results show that BridgeNet achieves significant performance improvements in multi-task learning, which were further validated through visual analysis and ablation experiments. Our work provides valuable insights and guidance for research and practice in the field of multi-task dense prediction. Future research can further explore the application and optimization of the BridgeNet method to meet the broader demands of dense prediction tasks.